use Euler’s method to obtain a four-decimal approximation of the indicated value. First use h = 0.1 and then use h = 0.05. Find an explicit solution for each initial-value problem


1
Expert's answer
2021-12-08T04:34:29-0500

"y'=2xy,y(1)=1,y(1.5)"


h = 0.1

"y_1=y_0+hf(x_0,y_0)=1+(0.1)f(1,1)=1.2"

"y_2=y_1+hf(x_1,y_1)=1.2+(0.1)f(1.1,1.2)=1.464"

"y_3=y_2+hf(x_2,y_2)=1.464+(0.1)f(1.2,1.464)=1.8154"

"y_4=y_3+hf(x_3,y_3)=1.8154+(0.1)f(1.3,1.8154)=2.2874"

 "y_5=y_4+hf(x_4,y_4)=2.2874+(0.1)f(1.4,2.2874)=2.9278"

"y(1.5)=2.9278"


h = 0.05

"y_1=y_0+hf(x_0,y_0)=1+(0.05)f(1,1)=1.1"

"y_2=y_1+hf(x_1,y_1)=1.1+(0.05)f(1.05,1.1)=1.2155"

"y_3=y_2+hf(x_2,y_2)=1.2155+(0.05)f(1.1,1.2155)=1.3492"

"y_4=y_3+hf(x_3,y_3)=1.3492+(0.05)f(1.15,1.3492)=1.5044"

"y_5=y_4+hf(x_4,y_4)=1.5044+(0.05)f(1.2,1.5044)=1.6849"

"y_6=y_5+hf(x_5,y_5)=1.6849+(0.05)f(1.25,1.6849)=1.8955"

"y_7=y_6+hf(x_6,y_6)=1.8955+(0.05)f(1.3,1.8955)=2.1419"

"y_8=y_7+hf(x_7,y_7)=2.1419+(0.05)f(1.35,2.1419)=2.4311"

"y_9=y_8+hf(x_8,y_8)=2.4311+(0.05)f(1.4,2.4311)=2.7714"

"y_{10}=y_9+hf(x_9,y_9)=2.7714+(0.05)f(1.45,2.7714)=3.1733"

"y(1.5)=3.1733"